### v1

@@ 6 List of Symbols
$c_{a}$ cost of action $a$, defined in Definition 2.1
$c_{a}^{k,\tau^{k}}$ cost functions of population $(k,\tau^{k})$, defined in (5.3)
$\boldsymbol{c}$ $(c_{a})_{a\in\mathcal{A}}$, defined in Definition 2.1
$\mathcal{I}$ $\mathcal{I}=\left(\boldsymbol{\gamma},\mathcal{T},\pi\right)$ , information structure, defined in Definition 2.3
@@
$p$ prior distribution on $\Theta$, defined in Definition 2.1
@@
$\mathcal{T}$ $\mathcal{T}=\times_{k\in\mathcal{K}}\mathcal{T}^{k}$, defined in Definition 2.3
$\mathcal{T}^{k}$ set of types for population $k$, defined in Definition 2.3
@@
$\boldsymbol{y}$ flow vector
$\boldsymbol{y}^{\ast}$ equilibrium flow vector
@@
$\mathcal{Y}$ $\Delta_{\gamma}(\mathcal{A})$, set of flow profiles, defined in Definition 2.1
$\gamma$ $\sum_{k\in\mathcal{K}}\gamma^{k}$ total population size, defined in Definition 2.1
$\gamma^{k}$ size of population $k$, defined in Definition 2.3
$\boldsymbol{\gamma}$ vector of population sizes, defined in Definition 2.3
$\Gamma$ $\left(\gamma,\mathcal{A},\Theta,p,\boldsymbol{c}\right)$, Bayesian nonatomic game, defined in Definition 2.1
@@
$\Delta(A)$ $\Delta_{1}(A)$, simplex of probability measures on $A$
@@
$\Theta$ state space, defined in Definition 2.1
$\mu$ outcome of the game, defined in Definition 2.1
@@
$\pi$ mapping from $\Theta$ to $\Delta(\mathcal{T})$, defined in Definition 2.3
@@
$\boldsymbol{\tau}=(\tau^{k})_{k\in\mathcal{K}}$ profile of types
$\Phi_{\theta}$ potential, defined in Definition 5.1

### v2

@@ Appendix A List of Symbols and Acronyms
$c_{e}$ cost of resource $e$
$C^{h}$ class of functions with a continuous $h$-th derivative
$e$ resource in a congestion game
$\mathcal{E}$ resource set in a congestion game
$\mathcal{I}$ $\mathcal{I}=\left(\boldsymbol{\gamma},\mathcal{T},\pi\right)$, information structure, defined in Definition˜2.3
@@
$L$ Lipschitz constant
$p$ probability distribution on $\Theta$, defined in Definition˜2.1
@@
$Q(\boldsymbol{y})$ aggregate quantity in an aggregative game, defined before (5.7)
$\mathcal{T}$ $\mathcal{T}=\times_{k\in\mathcal{K}}\mathcal{T}^{k}$, defined in Definition˜2.3
$\mathcal{T}^{k}$ set of types for population $k$, defined in Definition˜2.3
@@
$u_{a}$ payoff of action $a$, defined in Definition˜2.1
$u_{a}^{k,\tau^{k}}$ payoff functions of population $(k,\tau^{k})$, defined in (2.4)
$\boldsymbol{u}$ $(u_{a})_{a\in\mathcal{A}}$, defined in Definition˜2.1
$(u_{a}^{k,\tau^{k}}(\widehat{\boldsymbol{y}}))_{k,\tau^{k},a}$ interim payoff profile
$U_{a}^{k,\tau^{k}}$ payoff function of population $(k,\tau^{k})$, defined in (5.13)
$\operatorname{\mathsf{U}}(\boldsymbol{y},\theta)$ $\sum_{a}y_{a}u_{a}(\boldsymbol{y},\theta)$, defined in (5.28)
@@
$\boldsymbol{y}$ flow profile
$\boldsymbol{y}^{\ast}$ equilibrium flow profile
@@
$\mathcal{Y}$ $\Delta_{\gamma}(\mathcal{A})$, set of flow profiles, defined before Definition˜2.1
$\gamma^{k}$ size of population $k$, defined in Definition˜2.3
$\boldsymbol{\gamma}$ vector of population sizes, defined in Definition˜2.3
$\Gamma$ $\left(\mathcal{A},\Theta,p,\boldsymbol{u}\right)$, basic game, defined in Definition˜2.1
@@
$\Delta(A)$ $\Delta_{1}(A)$, simplex of probability measures on AA
@@
$\Theta$ state space, defined in Definition˜2.1
$\mu$ outcome of the game, defined after Definition˜2.1
@@
$\pi$ mapping from $\Theta$ to $\Delta(\mathcal{T})$, defined in Definition˜2.3
@@
$\boldsymbol{\tau}$ $(\tau^{k})_{k\in\mathcal{K}}$ type profile
$\Phi_{\theta}$ potential, defined in Definition˜5.1
$\psi$ designer welfare function, first used in Theorem˜4.1
@@
BCWE Bayes correlated Wardrop equilibrium, defined in Definition˜2.2
BWE Bayesian Wardrop equilibrium, defined after (2.6)
[…]